# Sylvester-Preconditioned Adaptive-Rank Implicit Time Integrators for Advection-Diffusion Equations with Variable Coefficients

### v2

@@ Abstract
We consider the adaptive-rank integration of 2D and 3D time-dependent advection-diffusion partial differential equations (PDEs) with variable coefficients. We employ a standard finite-difference method for spatial discretization coupled with diagonally implicit Runge-Kutta temporal schemes. The discrete equation is a generalized Sylvester equation (GSE), which we solve with an adaptive-rank algorithm structured around three key strategies: (i) constructing dimension-wise subspaces based on an extended Krylov strategy, (ii) developing an effective preconditioner for the reduced coefficient matrix, and (iii) efficiently computing the residual of the equation without explicitly reverting to the full-rank form. The low-rank decomposition is performed in 2D using SVD, and with high-order SVD (HOSVD) in 3D to represent the tensor in a compressed Tucker format. The computational complexity of the proposed approach is demonstrated numerically to be comparable to the constant-coefficient case [14], scaling as $\mathcal{O}(N{r^{2}}+{r^{d+1}})$ for $d$-dimensional problems (here, $d=2$ or $3$), with $N$ the resolution in one dimension and $r$ the maximal rank during the Krylov iteration (which we find to be largely independent of $N$). We present numerical examples that illustrate the computational efficacy and complexity of our algorithm.
@@
1 Introduction:
We propose an adaptive-rank numerical scheme for approximating solutions of time-dependent variable-coefficient advection-diffusion equations. The goal is to reduce the computational complexity of high-dimensional PDE solvers by exploiting the low-rank structure of the solution, building upon existing literature on low-rank PDE solvers [16, 13] and model order reduction (MOR) for solving high-dimensional linear and nonlinear systems [24, 16, 5, 40].
Two primary strategies in the literature exploit evolving low-rank structures for time-dependent PDEs: the dynamical low-rank (DLR) approximation [21, 27, 8, 11, 26] and the step-and-truncate (SAT) method, which has both explicit [23, 10, 18, 17, 19] and implicit [35, 30, 14, 28, 29] variants. Our study employs SAT, discretizing implicitly in time with high-order accuracy, leading to matrix/tensor differential equations to be solved at every time step. For constant coefficients in 2D, the advection-diffusion PDE becomes a Sylvester equation (SE) of the form:
@@ 1 Introduction:
where $\mathbf{X}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ represents the solution at time $t_{1}$, $\mathbf{X}_{0}\in\mathbb{R}^{N_{1}\times N_{2}}$ is the initial condition at $t_{0}$, and $A$ and $B$ are difference operators arising from the PDE discretization. Here, $N_{1}$ and $N_{2}$ are the 2D grid dimensions.
Krylov subspace methods have proven effective for solving SEs resulting from the discretization of PDEs [24, 39, 35, 31, 6]. In these methods, the solution bases are chosen from a Krylov (or Krylov-like) subspace of $A$ and $B$ in (1), and the original system is Galerkin-projected onto this reduced subspace. To accelerate the convergence of iterative methods, extensions of Krylov subspaces have been developed [12] involving $A^{-1}$, $B^{-1}$, which have been successfully applied to matrix equations [39, 14].
While substantial progress has been made for constant-coefficient problems, PDEs with variable coefficients remain challenging. Discretizing such systems leads to the so-called generalized Sylvester equation (GSE), of the form:
$\sum_{i=1}^{l}A_{i}\mathbf{X}_{1}{B}^{\top}_{i}=\mathbf{X}_{0},$ (2)
where $A_{i}$ and $B_{i}$ represent discretization operators for advection and/or diffusion terms. Several methods have been proposed to solve GSEs either with stationary iterative techniques  [5, 38] or Krylov methods  [5, 39, 32, 42]. Stationary iterative methods leverage low-rank techniques to enhance efficiency, achieving computational complexity of $\mathcal{O}(N)$ when implemented effectively [38]. However, their effectiveness critically depends on whether a dominating operator can be correctly identified to enable a contractive iterative process. For the Krylov approach, recent approaches leverage a Galerkin projection of the matrix equation on the dimension-wise Krylov bases when the right-hand side is low-rank [36, 40]. Compared to a standard Kronecker Krylov iteration, dimension-wise Krylov subspaces are inherently richer and yield superior approximations [32, 42, 24].
Several studies in the literature have considered using atypical extended Krylov subspaces to solve GSEs (see [5, 33, 20] and references therein). In [5], the authors present a series of techniques to tackle the Generalized Lyapunov equation, which is a special case of (2) of the form:
$A\mathbf{X}+\mathbf{X}A^{T}+\sum_{j=1}^{m}M_{j}\mathbf{X}M_{j}^{T}=\mathbf{F}.$ (3)
Among them, they discuss a low-rank projection-based scheme, in which extended Krylov subspaces are constructed using $A$ and standard Krylov subspaces are used for the remaining operators. In [20], a Krylov-based projection technique is developed for Generalized Lyapunov equations with $m=1$. The method builds its approximation subspaces by block augmentations generated from $A$ and $M_{1}$. In [33], a low-rank projection-based technique is employed for systems of disparate dimension sizes, $N_{2}\ll N_{1}$. The authors employ a nested sequence of block rational Krylov subspaces generated from all the operators in the $N_{1}$ direction, and solve the resulting small-sized equation with either matrix-based Conjugate Gradient or a Kronecker formulation (tractable since $N_{2}$ is small). In all of these studies, atypical approximation subspaces are formed by augmenting the basis with multiple applications of the operators in the multi‑term matrix equation, a strategy we adopt in this work.
Another Krylov-based line of inquiry, particularly relevant to our study, employs a basis-update-and-Galerkin (BUG) DLR integrator as a low-rank GMRES preconditioner for the GSE arising from anisotropic diffusion time-dependent problems with variable coefficients [28]. The BUG preconditioner evolves the basis and the coefficient matrix to provide guesses for the basis vectors and coefficient updates of the outer GMRES iteration, demonstrating significant potential in controlling GMRES iteration count and maximal Krylov rank. The approach is compatible with high-order temporal integrators. However, the reference offers no discussion of asymptotic computational complexity.
In this study, we propose to solve the GSE for time-dependent advection-diffusion equations leveraging adaptive-rank approximations of solutions on tensor-product grids with $N$ grid points per dimension, and using extended Krylov subspaces to update the basis vectors as well as the rank (denoted as $r$). Our main goal is to minimize the computational cost of the adaptive rank implementation without sacrificing accuracy. Key innovations of our proposed algorithm include:
  • Generalized Extended Krylov (xKrylov) Adaptive-Rank Integrators with Complexity Reduction for 2D and 3D Problems. We consider matrix solutions with factorizations of the form $X=USV^{\top}$, where basis vectors in $U$ and $V$ span the solution space and determine the rank of the solution. In the 2D setting, we employ a truncated singular value decomposition (SVD) to factorize the solution. In 3D, we use high-order singular value decomposition (HOSVD) to represent the 3D tensor in a compressed Tucker format. Our proposed approach constructs a set of xKrylov subspaces (for each dimension) that effectively capture the evolution of solution basis in an iterative manner. A Galerkin projection is then applied to project the full system to a reduced GSE. We propose a residual-checking mechanism that leverages the low-rank structure of solutions directly, thus avoiding forming the full residual matrix for norm evaluations. For the rest of the paper, we refer to the xKrylov iteration as the outer iteration.
  • Effective Preconditioning Strategy: The effectiveness of these projection-based approaches depends on the availability of an effective solver for the reduced GSE (2) of size up to $r$, the maximum size of dimension-wise xKrylov subspaces. There are two main challenges to solve it: (i) an increased computational load due to the multi-term nature of the equation, leading to rapid increase of xKrylov subspace sizes, and (ii) slow convergence of iterative methods (e.g., GMRES) resulting from algebraic ill-conditioning when solved iteratively. These challenges highlight the need for an effective preconditioner. Inspired by [31], we propose an averaged-coefficient Sylvester (ACS) approximation operator to the GSE as a preconditioner to GMRES. The approximate operator is of the Sylvester form and compact-equivalent to the original GSE, ensuring excellent performance as a GMRES accelerator [1, 2]. The direct inversion of the SE features a favorable computational complexity of $\mathcal{O}(r^{d+1})$, to be contrasted with the $\mathcal{O}(r^{3d})$ for $d$-dimensional problems when directly inverting the Kronecker-product form of the GSE. We refer to the preconditioned-GMRES reduced-GSE solve as the inner iteration.
  • Efficient Computational Complexity: As proposed, our method achieves a computational complexity of $\mathcal{O}(Nr^{2}+r^{d+1})$ for $d$-dimensional problems, offering a significant advantage over existing state-of-the-art adaptive-rank approaches. The $Nr^{2}$ term corresponds to the outer xKrylov iteration, and the $r^{d+1}$ term to the inner GMRES iteration, assuming an $r$-independent GMRES convergence rate (which we achieve with the ACS preconditioner). In principle, the rank $r$ may scale with $N$ [24], but we do not find this in practice. The favorable scaling properties of our approach will be verified in our numerical experiments.
@@
The rest of this paper is organized as follows. In Section 2, we consider the case of rank-one advection and diffusion coefficients and derive the GSE that results from a finite-difference discretization in space and backward Euler in time. In Section 3, we introduce an adaptive-rank solution strategy for 2D problems, including the coefficient-averaged approximation operator, construction of xKrylov subspaces, preconditioning technique in solving the reduced system, and low-rank computation of the residual norm. The extension of the technique to high-order temporal discretizations with multi-rank coefficients is provided in Appendix C, and for 3D problems in Appendix D. In Section 4, numerical results are presented to demonstrate the efficacy of the proposed algorithms. Finally, we conclude and provide outlook for future research in Section 5.
@@ 2 Problem Description and Generalized Sylvester Equation
For simplicity, we focus this section on 2D, with the extension to 3D elaborated in Appendix D. We establish the following notation conventions. Lowercase Greek letters, such as $\phi(x,y)$ or $\sigma(x,y)$, denote continuous scalar fields. Uppercase bold italic Greek letters, such as $\bm{\mathit{\Sigma}}(x,y)$ or $\bm{\mathit{\Phi}}(x,y)$, are used for vector-valued and tensor-valued functions. Uppercase Greek letters, like $\Phi$ or $\Sigma$, represent discrete matrices arising from the discretization of continuous scalar functions or differential operators on a grid. Boldface uppercase letters, such as $\mathbf{F}$ or $\mathbf{S}$, represent unknown matrix solutions or matrix right-hand sides in matrix equations. The identity matrix of size $n$ is denoted by ${I}_{n}$.
@@
$\sigma^{x}(x,y)=\sum_{i=1}^{k_{x}}\sigma^{1,x}_{i}(x)\sigma^{2,x}_{i}(y),\quad\sigma^{y}(x,y)=\sum_{i=1}^{k_{y}}\sigma^{1,y}_{i}(x)\sigma^{2,y}_{i}(y).$ (6)
@@
$\phi^{x}=\phi^{1,x}(x)\phi^{2,x}(y),\quad\phi^{y}=\phi^{1,y}(x)\phi^{2,y}(y),$
$\sigma^{x}=\sigma^{1,x}(x)\sigma^{2,x}(y),\quad\sigma^{y}=\sigma^{1,y}(x)\sigma^{2,y}(y).$
@@
$\Phi_{1}$ $=\text{diag}(\phi^{1,y}(x_{i})),\Sigma_{1}=\text{diag}(\sigma^{1,y}(x_{i})),\quad i=0,\cdots,N_{1}-1,$
$\Phi_{2}$ $=\text{diag}(\phi^{2,x}(y_{j})),\Sigma_{2}=\text{diag}(\sigma^{2,x}(y_{j})),\quad j=0,\cdots,N_{2}-1.$
@@
$T_{1}=D_{xx}^{\Phi},\quad T_{2}=D_{yy}^{\Phi},\quad T_{3}=-D_{x}^{\Sigma},\quad T_{4}=-D_{y}^{\Sigma}.$ (10)

### v3

@@ Abstract
We consider the adaptive-rank integration of multi-dimensional time-dependent advection-diffusion partial differential equations (PDEs) with variable coefficients. We employ a standard finite-difference method for spatial discretization coupled with high-order diagonally implicit Runge-Kutta temporal schemes. The discrete equation is a generalized Sylvester equation (GSE), which we solve with a projection-based adaptive-rank algorithm structured around two key strategies: (i) constructing dimension-wise subspaces using a novel atypical extended Krylov strategy, and (ii) efficiently solving the basis coefficient matrix with a preconditioned GMRES solver. The low-rank decomposition is performed in 2D using SVD and with high-order SVD (HOSVD) in 3D to represent the tensor in a compressed Tucker format. For $d$-dimensional problems (here, $d=2$ or $3$), the computational complexity and memory storage of the approach are found numerically to scale as $\mathcal{O}(N{r^{2}})+\mathcal{O}({r^{d+1}})$ and $\mathcal{O}(Nr)+\mathcal{O}(r^{d})$, respectively, with $N$ the one-dimensional resolution and $r$ the maximal rank during the Krylov iteration (which we find to be largely independent of $N$ on our numerical examples). We present numerical examples that illustrate the advertised properties of the algorithm.
@@
1 Introduction
We propose an adaptive-rank numerical scheme for approximating solutions of multi-dimensional time-dependent variable-coefficient advection-diffusion equations. The goal is to reduce the computational complexity of high-dimensional PDE solvers by dynamically exploiting the low-rank structure of the time-dependent solution when it exists, building upon existing literature on low-rank PDE solvers [16, 13] and model order reduction (MOR) for solving high-dimensional linear and nonlinear systems [25, 16, 5, 43].
Two primary strategies in the literature exploit evolving low-rank structures for time-dependent PDEs: the dynamical low-rank (DLR) approximation [22, 28, 8, 11] and the step-and-truncate (SAT) method, which has both explicit [24, 10, 18, 17, 19] and implicit [38, 31, 14, 29, 30] variants. Recently, [13] provided an overview of these two approaches with a focus on time-dependent kinetic simulations. Our study employs the SAT approach, which is known for preserving the high-order spatial and temporal accuracy of traditional mesh-based solvers for time-dependent problems. Specifically, to capture the solution dynamics accurately, we apply a high-order diagonally implicit Runge-Kutta (DIRK) time discretization to the matrix differential equation arising from the semi-discretization of advection-diffusion equations on a tensor-product of 1D spatial grids. The backward Euler time discretization of a 2D constant-coefficient advection-diffusion equation yields a Sylvester equation (SE) of the form:
@@ 1 Introduction
where $\mathbf{X}_{0},\mathbf{X}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ represent the solution snapshots at times $t_{0}$ and $t_{1}$, respectively, and $A$ and $B$ are coefficient matrices derived from the discretizations of advection and diffusion differential operators in two dimensions. Here $N_{1}$ and $N_{2}$ denote the number of grid points along each coordinate direction of the tensor product grid. Generalization to problems with variable advection and/or diffusion coefficients leads to the so-called generalized Sylvester equation (GSE) in matrix and Kronecker forms, respectively:
$\sum_{i=1}^{l}A_{i}\mathbf{X}_{1}{B}^{\top}_{i}=\mathbf{X}_{0}\qquad\iff\qquad\underbrace{\left(\sum_{i=1}^{l}{B}_{i}\otimes A_{i}\right)}_{\mathcal{A}}\text{vec}(\mathbf{X}_{1})=\text{vec}(\mathbf{X_{0}}),$ (2)
where $A_{i}$ and $B_{i}$ are coefficient matrices from discretization of the PDE operators, leveraging the separability of variable advection-diffusion coefficients. Higher-order DIRK schemes lead to SE and GSE in forms similar to (1)-(2), but with the right-hand side (RHS) term, $\mathbf{X}_{0}$, constructed from the previous DIRK stages. To exploit low-rank structure in matrices (for 2D) and tensors (for 3D and beyond), such RHS term can be efficiently compressed using SVD rounding in 2D, Tucker decomposition in 3D, and hierarchical Tucker or tensor train decomposition for problems beyond 3D [32, 27]. Ultimately, a successful development of efficient and robust solvers for the GSE (2) is of critical importance to enable adaptive-rank integrators that accurately capture the transient states of time-dependent problems.
Two broad families of algorithms are commonly employed to tackle GSEs (and SEs as a special case): projection-type methods [5, 20, 45, 36, 39, 25, 43, 41] and matrix-oriented classical Krylov methods [45, 35, 5, 34, 33, 26]. Projection-type schemes first construct low-dimensional subspaces that capture the dominant features of the solution. The GSE is then projected dimension-by-dimension via a Galerkin condition, yielding a much smaller GSE for the expansion coefficients that compose the bases together. Matrix-oriented Krylov iterations, in contrast, construct an approximation subspace by applying powers of the operator $\mathcal{A}$ in (2) to the vectorized RHS, but exploit the Kronecker structure so that the action of $\mathcal{A}$ is performed at the matrix level, and the full Kroneckerized operator is never formed explicitly. Inner products and other Krylov routines are likewise executed at the matrix-level, with low-rank truncation applied throughout to control memory. Although an appropriately chosen projection space can make a projection-type method formally equivalent to a matrix-oriented Krylov scheme, the projection-based framework offers additional freedom to design nonstandard subspaces that may accelerate convergence [45]. On the downside, projection methods are less amenable to precondition: their efficiency hinges on the efficacy of the chosen subspaces and on the accuracy with which the reduced problem is solved. Classical Krylov methods, in contrast, benefit from well-established preconditioning strategies but offer no flexibility to alter the search space, as the basis is typically generated by an Arnoldi or Lanczos-type process applied directly to $\mathcal{A}$. In summary, each approach has complementary strengths and weaknesses, and the choice between them should be guided by the structure of the particular GSE under investigation. Below we survey representative work in both directions.
In the context of solving SEs, projection-type methods based on Krylov subspace techniques have proven highly effective, particularly by exploiting the low-rank structure of PDE solutions [25, 42, 38, 6, 14]. These methods typically construct dimension-wise solution bases from Krylov or Krylov-like subspaces and apply a Galerkin projection onto the reduced subspace spanned by the outer product of dimension-wise Krylov subspaces [39, 25, 43]. To accelerate convergence, extended Krylov subspace methods (first proposed in [12] for the approximation of matrix functions) that incorporate $A$ and $A^{-1}$, as well as $B$ and $B^{-1}$, in each direction, respectively, were successfully applied to matrix equations in [42, 14] within a projection-based scheme termed Krylov-Plus-Inverted-Krylov (KPIK).
Applying projection-type methods to GSEs is much more challenging. The major difficulty lies in preventing the rapid growth of the solution approximation spaces due to the (possibly large) multi-term structure of the equation, which otherwise results in large reduced systems (of size $r^{d}$, where $r$ is the dimension of the per-direction Krylov subspace and $d$ is the number of dimensions). Such large reduced system demands efficient, tailored solution strategies. Generalized Lyapunov equations of the form ($A\mathbf{X}_{1}+\mathbf{X}_{1}A+\sum_{i=1}^{l}M_{i}X_{1}M_{i}=\mathbf{X}_{0}$), a special case of the GSE matrix equation (2), have received significant attention  [5, 20] exemplifying such challenges. In [5], projection spaces are built from extended Krylov subspaces from the operator $A$ and standard Krylov subspaces from the remaining operators ($M_{1},\cdots M_{l}$), but do not address how to solve the reduced system efficiently. In [20], for the case $l=1$, projection spaces are built from a rational Krylov from $A$ and inverted Krylov from $A+M_{1}$. The authors then use the Sherman-Woodbury formula to express the projected equation as a series of Lyapunov solves, relying on the availability of a low-rank decomposition of the Galerkin projected generalized operators in the system. The resulting complexity for the reduced system is $\mathcal{O}(r^{d+1})$ under the assumption of fast decaying singular values of the operators, but deteriorates to $\mathcal{O}(r^{3d})$ otherwise.
Low-rank matrix-oriented Krylov methods such as Conjugate Gradient (CG) or GMRES [5, 26, 33, 46] provide an alternative to projection-type solvers. In [5], several low-rank matrix-oriented Krylov methods, including CG and BiCGSTAB, are surveyed, and a comparative numerical study against projection-based methods is presented. The study shows that, across a few benchmark problems, projection methods (for the specific approximations spaces utilized) underperforms preconditioned BiCGSTAB Krylov method in terms of both convergence rate and memory usage. In contrast, Reference  [45] shows that projection-based methods can be computationally superior to the classical conjugate gradient method when an appropriate solution subspace is chosen, and that classical Krylov methods suffer from performance degradation due to loss of orthogonality from the low-rank truncation of the subspaces. Reference [34] confirms the loss of orthogonality issues in low-rank Krylov subspace methods, and proposes techniques to mitigate it. They further find that preconditioning is key to avoid substantial subspace growth in matrix-oriented low-rank Krylov methods.
Preconditioning is indeed key to low-rank Krylov solvers for GSEs. Reference [34] proposes an interesting preconditioner for GSEs stemming from variable advection–diffusion equations, which will be of particular relevance to the present study. By averaging the PDE coefficients—a strategy we term here the average-coefficient-Sylvester (ACS) approximation—they obtain a constant-coefficient preconditioner of an SE form. The GSE system is then solved using matrix-oriented GMRES (but without low-rank truncation), where the operator application exploits the Kronecker product structure of the GSE, and the preconditioner is applied using KPIK [42]. In another related work on anisotropic diffusion problems [29], a basis-update-and-Galerkin (BUG) DLR integrator is employed as a preconditioner for a low-rank matrix-oriented GMRES solver for the associated GSE. The BUG preconditioner evolves the basis and the coefficient matrix to provide effective initial guesses for both the basis vectors and coefficient updates in the GMRES iteration, demonstrating significant potential in reducing the GMRES iteration count and the maximal Krylov rank. Therefore, the performance of (low-rank) matrix-oriented Krylov methods critically depends on the availability of a cheap and suitable preconditioner and an efficient truncation mechanism to limit the iteration count and control subspace growth.
In this study, we build upon the advances above for SE and GSE solvers within the projection-based framework and develop an adaptive-rank implicit integrator for 2D and 3D time-dependent advection-diffusion equations with variable coefficients. The main innovation of the algorithm lies in the specific treatment of both the basis construction and coefficient-finding steps within a standard projection-based Krylov framework, in which we combine solution elements proposed elsewhere to produce a novel algorithm. Notably, from the projection-based framework, we borrow the idea of enriched atypical Krylov subspaces to ensure an efficient representation of the solution bases. From classical matrix-oriented Krylov methods, we adopt the idea of truncating the subspaces along the iterations to keep it as compact as possible, and the idea of preconditioning the GSE with a simpler equation that is spectrally close. We assemble these ingredients in the following way: For the basis-building step (which we will refer to as the outer iteration), we introduce a new Krylov subspace construction strategy that combines (i) inverted Krylov subspaces for each differentiation coefficient matrix, (ii) standard Krylov subspaces for diagonal matrices arising from separability of advection-diffusion coefficients, and (iii) extended Krylov subspaces for averaged-coefficient matrices formed from the ACS approximation. We control excessive growth of the rank of the so-constructed Krylov subspaces bases by a targeted truncation procedure using SVD-based QR and modified Gram-Schmidt orthogonalization. In the coefficient-finding step (which we refer to as the inner iteration), we follow a standard Galerkin projection and solve the associated system with GMRES, preconditioned by the SE resulting from the Galerkin projection of the ACS approximation. This preconditioner is expected to perform optimally because it is compact equivalent to the original system [1, 2] and therefore spectrally close. The key distinction of this approach from others in the literature is applying the ACS SE preconditioner to the reduced-system of the projection-based framework with $\mathcal{O}(r^{d+1})$, rather than to the full-scale system (as was done for instance in the matrix-oriented GMRES framework proposed in [35]).
These algorithmic innovations enable the method to achieve linear scaling with the one-dimensional resolution $N$, both for computational complexity [as $\mathcal{O}(Nr^{2})+\mathcal{O}(r^{d+1})$] and for memory storage [as $\mathcal{O}(Nr)+\mathcal{O}(r^{d})$], and allow us to perform 3D simulations of advection-diffusion problems with equivalent resolutions of $\mathcal{O}\left(10^{12}\right)$ degrees of freedom on a standard laptop, orders of magnitude faster than an optimally performing multigrid method scaling as $N^{d}\log(N)$ [21]. Additionally, as in [14], the adaptive low-rank integrator is amenable for use with DIRK methods for high-order temporal accuracy (up to third-order is considered in this study). All these advertised properties will be demonstrated numerically in the test examples.
@@
The rest of this paper is organized as follows. In Section 2, we consider the case of rank-one advection and diffusion coefficients and derive the GSE that results from a finite-difference discretization in space and backward Euler in time. In Section 3, we introduce an adaptive-rank solution strategy for 2D problems, including the coefficient-averaged approximation operator, the construction of extended Krylov subspaces (which we abbreviate as xKrylov in the sequel), the preconditioning technique in solving the reduced system, and the low-rank computation of the residual norm. The extension of the technique to high-order temporal discretizations with multi-rank coefficients is provided in Appendix C, and for 3D problems in Appendix D. In Section 4, numerical results are presented to demonstrate the efficacy of the algorithm. Finally, we conclude and provide an outlook for future research in Section 5.
@@ 2 Problem Description and Generalized Sylvester Equation
For simplicity, we focus the discussion in this section on 2D, with the extension to 3D elaborated in Appendix D. We establish the following notation conventions. Lowercase Greek letters, such as $\phi(x,y)$ or $\sigma(x,y)$, denote continuous scalar fields. Uppercase bold italic Greek letters, such as $\bm{\mathit{\Sigma}}(x,y)$ or $\bm{\mathit{\Phi}}(x,y)$, are used for vector-valued and tensor-valued functions. Uppercase Greek letters, like $\Phi$ or $\Sigma$, represent discrete matrices arising from the discretization of continuous scalar functions or differential operators on a grid. Boldface uppercase letters, such as $\mathbf{F}$ or $\mathbf{S}$, represent unknown matrix solutions or matrix right-hand sides in matrix equations. The identity matrix of size $n$ is denoted by ${I}_{n}$.
@@
$\sigma^{x}(x,y)=\sum_{i=1}^{k_{x}}\sigma^{1,x}_{i}(x)\sigma^{2,x}_{i}(y),\qquad\sigma^{y}(x,y)=\sum_{i=1}^{k_{y}}\sigma^{1,y}_{i}(x)\sigma^{2,y}_{i}(y).$ (5)
@@
$\phi^{x}=\phi^{1,x}(x)\phi^{2,x}(y),\qquad\phi^{y}=\phi^{1,y}(x)\phi^{2,y}(y),$
$\sigma^{x}=\sigma^{1,x}(x)\sigma^{2,x}(y),\qquad\sigma^{y}=\sigma^{1,y}(x)\sigma^{2,y}(y).$
@@
$\Phi_{1}$ $=\text{diag}(\phi^{1,y}(x_{i})),\Sigma_{1}=\text{diag}(\sigma^{1,y}(x_{i})),\qquad i=0,\cdots,N_{1}-1,$
$\Phi_{2}$ $=\text{diag}(\phi^{2,x}(y_{j})),\Sigma_{2}=\text{diag}(\sigma^{2,x}(y_{j})),\qquad j=0,\cdots,N_{2}-1.$
@@
$T_{1}=D_{xx}^{\Phi},\qquad T_{2}=D_{yy}^{\Phi},\qquad T_{3}=-D_{x}^{\Sigma},\qquad T_{4}=-D_{y}^{\Sigma}.$ (9)